OpenStudy (anonymous):

How do you get the answer for 10/12 - 2/3 ? Please include steps.

OpenStudy (anonymous):

10/2=5 2/3=about 0.66 So 5-0.66=about 4.34. Hope this helps.

OpenStudy (anonymous):

But the answer is 1/6. Idk how to get that

OpenStudy (anonymous):

@lunatiger

OpenStudy (anonymous):

If it is fractions it would be 10/12 you divide by the gcf and get 5/6 then you multiply 2/3 by 2 and get 4/6 5/6-4/6=1/6

OpenStudy (anonymous):

Thank you. @lunatiger

OpenStudy (anonymous):

You're welcome.